Ok, in a bit of a pickle. I was given the following: F(x)= |x-2|
From here I am given a series of about 10 different transformations I am supposed to make using the F(x) provided and these transformations are things including shifting left and right, reflections, or a combination of changes. I have the original graphed and the coordinates I have are (-2,4), (-1,3), (0,2), (1,1), (2,0), (3,1), (4,2), (5,3). Then I was given the following transformation that I am supposed to apply to the original: f(x-2). Do I set it up as f(x-2) = |x-2| and get new x and y coordinates? Or do I just change the x value and keep the y's the same? When I change the x's and y's I end up with the original coordinates I have for the original function, so I assume I only change the x and keep the y's the same? Please help!!! I am also given other changes I am supposed to apply to the original function, which again is: F(x)=|x-2| The changes I am given are things such as : f(x+4), f(x+1)-2 , 3f(x) and f(2x) . What are these asking me to do?!?!

Answer by solver91311(24713)



If F(x) = x - 2, then F(1) = 1 - 2, F(a) = a - 2, and F(x - 2) = (x - 2) - 2.

See how far you get applying that.

John

Egw to Beta kai to Sigma
My calculator said it, I believe it, that settles it
The Out Campaign: Scarlet Letter of Atheism